I need the work for the problem x/2-7=15
Answer by lynnlo(4176)

X/2-7=15
X/2-7-15=0==MOVE -15 TO THE LEFT
X/2+-22=0==ADD FRACTIONS OR INTEGERS TOGETHER
X/2-22=0==REMOVE THE + SIGN IN FRONT OF THE -22
SOLVE THE LINEAR EQUATION
X/2-22=0
THIS IS EQUIVALENT TO 0.5*X-22-0
RESULT========0=0
SOLUTION======X=44